he Nelson Company has $1,625,000 in current assets and $650,000 in current liabilities. Its initial inventory level is $520,000, and it will raise funds as additional notes payable and use them to increase inventory.

  1. How much can Nelson's short-term debt (notes payable) increase without pushing its current ratio below 1.5? Round your answer to the nearest cent.

    $  


  2. What will be the firm's quick ratio after Nelson has raised the maximum amount of short-term funds? Round your answer to two decimal places.
Answer #1
Req 1.
Current Assets:
Inventory 520000
Liquid assets 1105000
(1625000-520000)
Total Current Assets 1625000
Current Liabilities: 650000
Let, short term debt by 'x'
Therefore,
Revised Current assets = 1625000+x
Revised Current Liabilities= 650000 +x
Current ratio shall be atleast 1.50 (Current ratio = Current Assets/ Current Liabilities)
Therefore, Maximum Short term debt shall be computed by following equation:
1.50 = 1625000 +x / 650000+x
975000+ 1.50 x= 1625000 + x
1.50x- x = 650000
0.50x = 650000
x= $ 1300,000
Therefore, Increase in Short term debt= $ 1300,000
Req 2.
Revised Current liabilities = 650000+1300000= 1950000
Liquid assets (Remain same) = 1105000
Revised Quick ratio = Liquid assets / Revised Current liabilities
1105000 /1950000 = 0.57